OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5? y = 5x − 3 y = 5x − 1 y = 5x + 3 y = 5x − 19

OpenStudy (uri):



OpenStudy (uri):

Formula is:y-y1=m(x-x1)

OpenStudy (uri):

Can you do it now? @jpw

OpenStudy (anonymous):

@jpw You want the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aWhat is the equation of the line that passes through the points (−1, 3) and (1, 11)? 4x − y = 1 4x + y = 1 x − 4y = −13 4x − y = −7 nd can u do this one to

OpenStudy (anonymous):

Number one is A

OpenStudy (anonymous):

what about the second one

OpenStudy (uri):

@DingDongLingLong We are not here to give answers..but guide users to to the answer..Please don't give answers again.Thanks. You need to look at it: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

He/she doesnt want the process. They asked for the answer i gave it. @uri

OpenStudy (uri):

So if anyone tells you to jump of a hill,will you? If someone is breaking the Coc will you too? @DingDongLingLong

OpenStudy (anonymous):

:P i see. And well, depends whats at the bottom of a hill. @uri

OpenStudy (anonymous):

can u do the second one